OpenStudy (anonymous):

How many atoms of phosphorus are in 5.20 mol of copper(II) phosphate?

OpenStudy (anonymous):

\[5.2 mol Cu _{3}(PO _{4})_{2}\left( \frac{2 mol P}{1 mol Cu _{3}(PO _{4})_{2}} \right)\left( \frac{ 6.022\times10^{23} atom}{ 1 mol P} \right)=6.26\times10^{24}\]